# Unlocking State-Tracking in Linear RNNs Through Negative Eigenvalues

### v3

@@ 3.2 Formal Language Theory
One LRNN Layer is an automaton.  Given an alphabet $\Sigma\subset\mathbb{N}$, we can view one layer of an LRNN in (1) as the automaton $\mathcal{A}_{\mathrm{lin}}=(\Sigma,\mathcal{H},{\bm{H}}_{0},\delta_{\mathrm{lin}})$, where $\delta_{\mathrm{lin}}({\bm{H}},w)={\bm{A}}(w){\bm{H}}+{\bm{B}}(w)$, which is extended as we saw previously [^1] We let $\delta_{\mathrm{lin}}:\mathbb{R}^{n{\times}d}\,{\times}\,\Sigma\,{\to}\,\mathbb{R}^{n\times d}$ and extend it to $\delta_{\mathrm{lin}}:\mathbb{R}^{n{\times}d}\,{\times}\,\Sigma^{*}\,{\to}\,\mathbb{R}^{n\times d}$, then we define $\mathcal{H}$., and =$\mathcal{H}=\{\delta_{\mathrm{lin}}({\bm{H}}_{0},{\bm{w}})\,:\,{\bm{w}}\in\Sigma^{*}\}\subseteq\mathbb{R}^{n\times d}$. We say that an LRNN layer in (1) implements the FSA $\mathcal{A}=(\Sigma,Q,q_{0},\delta)$ if $\mathcal{A}_{\mathrm{lin}}$ can mimic the state transitions of $\mathcal{A}$ [^2] This definition is equivalent to that of FSA homomorphism, see (Maler & Pnueli, 1994, Definition 3).. Formally, if there exists a surjective function $g:\mathcal{H}\to Q$, such that for any ${\bm{H}}\in\mathcal{H}$, $w\in\Sigma$
$\quad\delta(g({\bm{H}}),w)=g(\delta_{\mathrm{lin}}({\bm{H}},w))=g({\bm{A}}(w){\bm{H}}+{\bm{B}}(w))$
Every language $L$ recognized by $\mathcal{A}$ can also be recognized by this LRNN layer with a sufficiently powerful $\mathrm{dec}$. In particular if $R\,{\subseteq}\,Q$ is the recognizing set for $L$ and $q_{0}=g({\bm{H}}_{0})$, then the decoder $\mathrm{dec}({\bm{H}}_{t},w_{t})=\mathbf{1}\{g({\bm{H}}_{t})\in R\}$, will correctly determine if ${\bm{w}}\in L$. Therefore, implementing $\mathcal{A}$ is at least as hard as recognizing $L$. A principal goal of this work is to show that current LRNNs cannot recognize simple languages such as parity (negative results) while appropriate modifications to the state-transition matrices, enable LRNNs to implement broader classes of FSA (positive results), with certain classes of FSA requiring a single layer. Note, that while LRNNs with one layer can recognize any regular language, the state transition matrices might not fit into the structure imposed by current LRNNs, such as those in Table 1 (see Section A.3 for more details).
@@ 4 Theoretical Analysis
We begin by highlighting the limitations of current LRNNs, demonstrating that they fail to meet a necessary condition for solving parity and modular counting problems: the eigenvalues of their state-transition matrices are restricted to the range $[0,1]$. Subsequently, we illustrate how extending this eigenvalue range to $[-1,1]$ significantly enhances the expressive power of LRNNs.
@@ 4.1 Limitations of Current LRNNs
The parity $y_{t}\in\{0,1\}$ of a sequence of ones and zeros $x_{1}\dots x_{t}\in\{0,1\}^{t}$ is 1 if the total number of ones in the sequence is odd, and 0 if it’s even. Equivalent to addition modulo 2, it can be computed by summing the values in the input sequence and then applying the modulo 2 function: $y_{t}=(\sum_{i=1}^{t}x_{i})\bmod 2$. We can also express this as the linear recurrence
$h_{t}=h_{t-1}+x_{t},\quad h_{0}=0,\quad y_{t}=h_{t}\bmod 2$ (2)
where $h_{t}$ contains the total number of ones. This solution can be implemented by an LRNN with one layer and scalar states by setting ${\bm{A}}(x_{t})=1$, ${\bm{B}}(x_{t})=x_{t}$, ${\bm{H}}_{0}=0$, and $\mathrm{dec}({\bm{H}}_{t},x_{t})={\bm{H}}_{t}\bmod 2$ in (1). However, implementing such a solution with finite precision presents an issue: the state $h_{t}$ can grow indefinitely with $t$, eventually reaching the limit of our precision range. Indeed, $h_{t}\in\{0,\dots,t\}$, requiring $\log_{2}(t+1)$ bits for storage. Moreover, in practice $\mathrm{dec}$ must approximate the modulus 2 function, which is challenging to learn due to its discontinuous and periodic nature. Such solutions, referred to as shortcut solutions, are the only ones learnable by Transformers when allowing $O(\log(t))$ bits of precision and either depth $O(\log(t))$ or width $O(t)$ (Liu et al., 2023).
A more efficient solution, which implements the two-state FSA solving this problem, can still be realized by a finite precision LRNN with one layer and scalar states (and consequently also with vector states and diagonal state-transition matrices) using the recurrence
$h_{t}=a(x_{t})h_{t-1}+b(x_{t}),\quad h_{0}=b(0)=0,\quad b(1)=a(0)=1,\quad{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a(1)=-1},\quad y_{t}=h_{t}.$ (3)
Note that the state-transition scalar $a(1)$ is negative, while current diagonal LRNNs do not allow negative values. (Sarrof et al., 2024, Theorem 2) states that this fact makes real-valued diagonal LRNNs unable to solve parity, which raises the question: can non-diagonal LRNNs which allow only positive eigenvalues, such as DeltaNet, solve parity? The following result answers this question negatively by generalizing Sarrof et al. (2024, Theorem 2) to non-diagonal matrices. To solve parity, the state transition matrices must allow at least one eigenvalue to be neither real nor positive. For non-diagonal matrices, this eigenvalue could simply have nonzero imaginary part.
@@ 4.2 Allowing Negative Eigenvalues
LRNNs with the modified state transition matrices can implement the solution to parity in (3) by setting $s(1)=0$ and $\phi(1)=1$ so that if we consider a scalar recursion, then ${\bm{A}}^{-}_{\mathrm{diag}}(1)=-1$. However, Theorem 2 shows that we cannot count modulo $3$ with diagonal state transition matrices, even when allowing negative eigenvalues. Despite this, it is well known that counting modulo $m$ can be achieved by rotating a vector in $\mathbb{R}^{2}$ by an angle of $2\pi/m$ radians, and we can express a rotation matrix as a product of two reflection matrices, which are GH matrices with eigenvalues in $\{-1,1\}$. In other words, for any $m\in\mathbb{N}$ there exist unit norm vectors ${\bm{v}}_{1},{\bm{v}}_{2}\in\mathbb{R}^{2}$ such that
Therefore, in the next section, we examine the impact of our change to the eigenvalue range on state-transition matrices constructed as repeated products of GH matrices.
@@ 4.3 Expressivity of Products of Generalized Householder Matrices
For any $n,k\in\mathbb{N}$, we define the set of all matrices in $\mathbb{R}^{n\times n}$ that can be expressed as a product of $k$ GH matrices, each having the only interesting eigenvalue in the range $\Omega\subseteq\mathbb{R}$, as
@@
Intuitively, higher $k$ means higher expressivity but also higher cost for matrix-vector products. Furthermore, as long as $\Omega\subseteq[-1,1]$, the norm of the matrices is bounded by one. We observe that if ${\bm{M}}\in\mathcal{M}_{1}^{n}(\{-1\})$, then ${\bm{M}}$ is a reflection (or Householder) matrix, and that for any ${\bm{x}}\in\mathbb{R}^{l}$, ${\bm{A}}_{\mathrm{GH}}({\bm{x}})\in\mathcal{M}_{1}^{n}([0,1])$ and ${\bm{A}}^{-}_{\mathrm{GH}}({\bm{x}})\in\mathcal{M}_{1}^{n}([-1,1])$ so that with our change we also include reflections. Moreover, $\mathcal{M}_{k}^{n}(\Omega)\subseteq\mathcal{M}_{k^{\prime}}^{n}(\Omega^{\prime})$ if $\Omega\subseteq\Omega^{\prime}$ and either $k^{\prime}=k$ or $k^{\prime}\geq k$, $1\in\Omega$.
@@
In the proof in Section C.2, we map each state-transition function to a matrix representation. This can always be done using permutation matrices, but for cyclic groups, we can also use rotation matrices. In the case of permutations, if every state-transition permutes at most $k$ states then the corresponding permutation matrix will be in $\mathcal{M}_{k-1}^{n}(\{-1,1\})$, since it is either the identity or can be written as a product of at most $k-1$ permutations of two elements (swaps), each in $\mathcal{M}_{1}^{n}(\{-1\})$.
A consequence of Theorem 3 is that if every transition function of the FSA has a permutation representation corresponding to a swap or the identity, then an LRNN layer with ${\bm{A}}={\bm{A}}^{-}_{\mathrm{GH}}$, can implement it. This is useful in practice because the time complexity of the LRNN having a product of $k$ GH matrices as one state-transition matrix increases linearly with $k$. Also, for natural language tasks, the state-transitions for the FSA might be either simple or encoded using multiple letters. For example, for addition modulo $5$, a word may look like “3+2+4=4” (two letters per addition). This allows an LRNN with state-transition matrices in $\mathcal{M}_{1}^{n}([-1,1])$ to model complex transitions. Indeed, if each transition uses $k$ letters and we set ${\bm{B}}\equiv 0$ and ${\bm{A}}:\mathbb{R}^{l}\to\mathcal{M}_{1}^{n}([-1,1])$ in (1), then the LRNN layer can model permutations that change up to $k+1$ elements since
@@
The proof in Section C.4 exploits the landmark Theorem by Krohn & Rhodes (1965), which states that every FSA can be decomposed as a cascade of simpler FSAs whose state-transition functions are either one-to-one or constant. Each layer of the LRNN will implement one FSA (with $n$ states) of the cascade using $n\times n$ permutation matrices, which are in $\mathcal{M}_{n-1}^{n}(\{-1,1\})$, for the one-to-one transitions, while for constant (state-independent) transitions it will set the corresponding state-transition matrix to $0\in\mathcal{M}_{n}^{n}(\{0\})$ and the function ${\bm{B}}$ appropriately. Note, that we can obtain the zero matrix only inefficiently as a product of $n$ GH matrices, while it could also be obtained with a single diagonal matrix. This points towards LRNNs using a mix of GH and diagonal matrices, as recently explored by Gated DeltaNet (Yang et al., 2025) and RWKV-7.
@@ 5.1 Chomsky Hierarchy
Transformer 0.022 0.031 0.025
mLSTM 0.087 (0.04) 0.040 (0.04) 0.034 (0.03)
sLSTM 1.000 (1.00) 0.787 (1.00) 0.173 (0.57)
Mamba $[0,1]$ 0.000 0.095 0.092
Mamba $[-1,1]$ 1.000 0.241 0.136
DeltaNet $[0,1]$ 0.017 0.314 0.137
DeltaNet $[-1,1]$ 1.000 0.971 0.200
We conducted experiments with some of the formal language tasks proposed by Deletang et al. (2023) and similarly used to benchmark xLSTM (Beck et al., 2024). Our focus was on tasks where mLSTM (an LRNN) previously underperformed while sLSTM (a non-linear RNN) succeeded, specifically parity, modular arithmetic without brackets (both regular languages). and modular arithmetic with brackets (context-free language). As in Beck et al. (2024), we trained each model with sequence lengths ranging from 3 to 40 and evaluated on lengths from 40 to 256, to assess length generalization. Note that our theoretical results cover just regular languages, excluding modular arithmetic with brackets. We compared a Transformer, mLSTM and sLSTM against two variants each of Mamba and DeltaNet - with and without eigenvalue range extension.
@@ Appendix C Products of Generalized Householder Matrices – Proofs
We provide proofs for the results stated in Section 4.3.
C.1 Proof of Proposition 1

### v5

@@ 3.2 Formal Language Theory
One LRNN Layer is an automaton.  Given an alphabet $\Sigma\subset\mathbb{N}$, we can view one layer of an LRNN in (1) as the automaton $\mathcal{A}_{\mathrm{lin}}=(\Sigma,\mathcal{H},{\bm{H}}_{0},\delta_{\mathrm{lin}})$, where $\delta_{\mathrm{lin}}({\bm{H}},w)={\bm{A}}(w){\bm{H}}+{\bm{B}}(w)$, which is extended as we saw previously [^1] We let $\delta_{\mathrm{lin}}:\mathbb{R}^{n{\times}d}\,{\times}\,\Sigma\,{\to}\,\mathbb{R}^{n\times d}$ and extend it to $\delta_{\mathrm{lin}}:\mathbb{R}^{n{\times}d}\,{\times}\,\Sigma^{*}\,{\to}\,\mathbb{R}^{n\times d}$, then we define $\mathcal{H}$., and =$\mathcal{H}=\{\delta_{\mathrm{lin}}({\bm{H}}_{0},{\bm{w}})\,:\,{\bm{w}}\in\Sigma^{*}\}\subseteq\mathbb{R}^{n\times d}$. We say that an LRNN layer in (1) implements the FSA $\mathcal{A}=(\Sigma,Q,q_{0},\delta)$ if $\mathcal{A}_{\mathrm{lin}}$ can mimic the state transitions of $\mathcal{A}$ [^2] This definition is equivalent to that of FSA homomorphism, see (Maler & Pnueli, 1994, Definition 3).. Formally, if there exists a surjective function $g:\mathcal{H}\to Q$, such that for any ${\bm{H}}\in\mathcal{H}$, $w\in\Sigma$ $\quad\delta(g({\bm{H}}),w)=g(\delta_{\mathrm{lin}}({\bm{H}},w))=g({\bm{A}}(w){\bm{H}}+{\bm{B}}(w))$. Every language $L$ recognized by $\mathcal{A}$ can also be recognized by this LRNN layer with a sufficiently powerful $\mathrm{dec}$. In particular if $R\,{\subseteq}\,Q$ is the recognizing set for $L$ and $q_{0}=g({\bm{H}}_{0})$, then the decoder $\mathrm{dec}({\bm{H}}_{t},w_{t})=\mathbf{1}\{g({\bm{H}}_{t})\in R\}$, will correctly determine if ${\bm{w}}\in L$. Therefore, implementing $\mathcal{A}$ is at least as hard as recognizing $L$. A principal goal of this work is to show that current LRNNs cannot recognize simple languages such as parity (negative results) while appropriate modifications to the state-transition matrices, enable LRNNs to implement broader classes of FSA (positive results), with certain classes of FSA requiring a single layer. Note, that while LRNNs with one layer can recognize any regular language, the state transition matrices might not fit into the structure imposed by current LRNNs, such as those in Table 1 (see Section A.3 for more details).
@@ 4.1 Limitations of Current LRNNs
t States converge or diverge ${\bm{h}}_{t}={\bm{A}}(1){\bm{h}}_{t-1}+{\bm{B}}(1)$, $\ \mathrm{eigs}({\bm{A}}(1))\geq 0$tOddEven Clear state separation $h_{t}=a(1)h_{t+1}+b(1),\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a(1)=-1}$$\mathbf{Even}$$\mathbf{Odd}$$11$Parity automaton
Figure 2: Parity requires negative eigenvalues. States of one-layer LRNNs with the sequence $1111\ldots$ as input. If the eigenvalues of $\mathbf{A}(1)$ are nonnegative, the states either diverge or converge monotonically, and so, for large enough $t$ and in finite precision, cannot be distinguished. In contrast, the LRNN with $a(1)=-1$ alternates between two states like the parity automaton.
In this section, we describe how positive eigenvalues and non-triangular state transition matrices limit LRNNs state-tracking capabitlies. In particular, we focus on parity and modular addition. The parity $y_{t}\in\{0,1\}$ of a sequence of ones and zeros $x_{1}\dots x_{t}\in\{0,1\}^{t}$ is 1 if the total number of ones in the sequence is odd, and 0 if it’s even. Equivalent to addition modulo 2, it can be computed by summing the values in the input sequence and then applying the modulo 2 function: $y_{t}=(\sum_{i=1}^{t}x_{i})\bmod 2$. This solution can be implemented by an LRNN with one layer and scalar states by setting ${\bm{A}}(x_{t})=1$, ${\bm{B}}(x_{t})=x_{t}$, ${\bm{H}}_{0}=0$, and $\mathrm{dec}({\bm{H}}_{t},x_{t})={\bm{H}}_{t}\bmod 2$ in (1). However, implementing such a solution with finite precision presents an issue: the state $h_{t}$ can grow indefinitely with $t$, eventually reaching the limit of our precision range. Indeed, $h_{t}\in\{0,\dots,t\}$, requiring $\log_{2}(t+1)$ bits for storage. Moreover, in practice $\mathrm{dec}$ must approximate the modulus 2 function, which is challenging to learn due to its discontinuous and periodic nature.
A more efficient solution, which implements the two-state FSA solving this problem, can still be realized by a finite precision LRNN with one layer and scalar states (and consequently also with vector states and diagonal state-transition matrices) using the recurrence $h_{t}=a(x_{t})h_{t-1}+b(x_{t}),\quad h_{0}=b(0)=0,\quad b(1)=a(0)=1,\quad{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a(1)=-1},\quad y_{t}=h_{t}$. Note that the state-transition scalar $a(1)$ is negative, while current diagonal LRNNs do not allow negative values. (Sarrof et al., 2024, Theorem 2) states that this fact makes real-valued diagonal LRNNs unable to solve parity, which raises the question: can non-diagonal LRNNs which allow only positive eigenvalues, such as DeltaNet, solve parity? The following result answers this question negatively by generalizing Sarrof et al. (2024, Theorem 2) to non-diagonal matrices. To solve parity, the state transition matrices must allow at least one eigenvalue to be neither real nor positive. For non-diagonal matrices, this eigenvalue could simply have nonzero imaginary part. The main idea of the theorem is illustrated in Figure 2.
@@ 4.2 Allowing Negative Eigenvalues
LRNNs with the modified state transition matrices can implement the solution to parity in (2) by setting $s(1)=0$ and $\phi(1)=1$ so that if we consider a scalar recursion, then ${\bm{A}}^{-}_{\mathrm{diag}}(1)=-1$. However, Theorem 2 shows that we cannot count modulo $3$ with triangular state transition matrices, even when allowing negative eigenvalues. Therefore, in the next section, we examine the impact of our change to the eigenvalue range on non-triangular state-transition matrices.
@@ 4.3 Expressivity of Products of Generalized Householder Matrices
We focus on state-transition matrices that are products of $k$ GH matrices. For DeltaNet $k=1$. For any $n,k\in\mathbb{N}$, we define the set of all matrices in $\mathbb{R}^{n\times n}$ that can be expressed as a product of $k$ GH matrices, each having the only interesting eigenvalue in the range $\Omega\subseteq\mathbb{R}$, as
@@
Intuitively, higher $k$ means higher expressivity but also higher cost for matrix-vector products. Furthermore, as long as $\Omega\subseteq[-1,1]$, the norm of the matrices is bounded by one, which guarantees that repeated matrix product do not diverge. We observe that if ${\bm{M}}\in\mathcal{M}_{1}^{n}(\{-1\})$, then ${\bm{M}}$ is a reflection (or Householder) matrix, and that for any ${\bm{x}}\in\mathbb{R}^{l}$, ${\bm{A}}_{\mathrm{GH}}({\bm{x}})\in\mathcal{M}_{1}^{n}([0,1])$ and ${\bm{A}}^{-}_{\mathrm{GH}}({\bm{x}})\in\mathcal{M}_{1}^{n}([-1,1])$ so that with our change we also include reflections. Moreover, $\mathcal{M}_{k}^{n}(\Omega)\subseteq\mathcal{M}_{k^{\prime}}^{n}(\Omega^{\prime})$ if $\Omega\subseteq\Omega^{\prime}$ and either $k^{\prime}=k$ or $k^{\prime}\geq k$, $1\in\Omega$.
@@
213132231swapswap$\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&1\end{pmatrix}$${\bm{I}}-2{\bm{v}}_{1}{\bm{v}}_{1}^{\top}$$\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix}$${\bm{I}}-2{\bm{v}}_{2}{\bm{v}}_{2}^{\top}$$\begin{pmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{pmatrix}$$\times$$=$
Figure 3: A permutation of $k$ elements is also a composition of at most $k{-}1$ swaps. This maps to a product of $k{-}1$ Hoseholders, each representing a swap. Illustrated for $k=3$.
$\mathbf{v}_{1}^{\top}{=}\left(\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}},0\right)$, $\mathbf{v}_{2}^{\top}{=}\left(0,\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}}\right)$.
@@
In the proof in Section C.3, we map each state-transition function to a matrix representation. This can always be done using permutation matrices, but for cyclic groups, we can also use rotation matrices (Section C.1). For permutations, if every state-transition permutes at most $k$ states then the corresponding permutation matrix will be in $\mathcal{M}_{k-1}^{n}(\{-1,1\})$, since it is either the identity or can be written as a product of at most $k-1$ permutations of two elements (swaps), each in $\mathcal{M}_{1}^{n}(\{-1\})$ (see Figure 3). A consequence of Theorem 3 is that if every transition function of the FSA has a permutation representation corresponding to a swap or the identity, then an LRNN layer with ${\bm{A}}={\bm{A}}^{-}_{\mathrm{GH}}$, can implement it. This is useful in practice because the time complexity of an LRNN having a product of $k$ GH matrices as one state-transition matrix increases linearly with $k$. Also, for natural language tasks, the state-transitions for the FSA might be either simple or encoded using multiple letters. For example, for addition modulo $5$, a word may look like “3+2+4=4” (two letters per addition). This allows an LRNN with state-transition matrices in $\mathcal{M}_{1}^{n}([-1,1])$ to model complex transitions. Indeed, if each transition uses $k$ letters and we set ${\bm{B}}\equiv 0$ and ${\bm{A}}:\mathbb{R}^{l}\to\mathcal{M}_{1}^{n}([-1,1])$ in (1), then the LRNN layer can model permutations that change up to $k+1$ elements since
@@
The proof in Section C.5 exploits the landmark Theorem by Krohn & Rhodes (1965), which states that every FSA can be decomposed as a cascade of simpler FSAs whose state-transition functions are either one-to-one or constant. Each layer of the LRNN will implement one FSA (with $n$ states) of the cascade using $n\times n$ permutation matrices, which are in $\mathcal{M}_{n-1}^{n}(\{-1,1\})$, for the one-to-one transitions, while for constant (state-independent) transitions it will set the corresponding state-transition matrix to $0\in\mathcal{M}_{n}^{n}(\{0\})$ and the function ${\bm{B}}$ appropriately. Note that we can obtain the zero matrix only inefficiently as a product of $n$ GH matrices, while it could also be obtained with a single diagonal matrix. This points towards LRNNs using a mix of GH and diagonal matrices, as recently explored by Gated DeltaNet (Yang et al., 2025) and RWKV-7.
@@ 5.1 Chomsky Hierarchy
Transformer 0.022 0.031 0.067
mLSTM 0.087 (0.04) 0.040 (0.04) 0.114 (0.03)
sLSTM 1.000 (1.00) 0.787 (1.00) 0.178 (0.57)
Mamba $[0,1]$ 0.000 0.095 0.123
Mamba $[-1,1]$ 1.000 0.241 0.116
DeltaNet $[0,1]$ 0.017 0.314 0.194
DeltaNet $[-1,1]$ 1.000 0.971 0.260
We conducted experiments with some of the formal language tasks proposed by Deletang et al. (2023) and similarly used to benchmark xLSTM (Beck et al., 2024). Our focus was on tasks where mLSTM (an LRNN) previously underperformed while sLSTM (a non-linear RNN) succeeded, specifically parity, modular arithmetic without brackets (both regular languages) and modular arithmetic with brackets (context-free language). As in Beck et al. (2024), we trained each model with sequence lengths ranging from 3 to 40 and evaluated on lengths from 40 to 256, to assess length generalization. Note that our theoretical results cover just regular languages, excluding modular arithmetic with brackets. We compared a Transformer, mLSTM and sLSTM against two variants each of Mamba and DeltaNet - with and without eigenvalue range extension.
@@ Appendix C Products of Generalized Householder Matrices – Proofs
We provide proofs for the results stated in Section 4.3. Before that, we illustrate how a linear RNN with one layer and state transition matrices that are products of 2 Householder matrices can count modulo $m$.
C.1 Products of Two Householders and Modular Counting
Counting modulo $m$ can be achieved by rotating a vector in $\mathbb{R}^{2}$ by an angle of $2\pi/m$ radians, and we can express a rotation matrix as a product of two reflection matrices, which are GH matrices with eigenvalues in $\{-1,1\}$ (see Section C.1). Inded, for any $m\in\mathbb{N}$ there exist unit norm vectors ${\bm{v}}_{1},{\bm{v}}_{2}\in\mathbb{R}^{2}$ such that
C.2 Proof of Proposition 1